OpenStudy (anonymous):

2.5x+1.75=-3.25+1.5x

OpenStudy (anonymous):

\(\large 2.5x+1.75=-3.25+1.5x\)\(\large 2.5x-1.5x+1.75-1.75=-3.25-1.75+1.5x-1.5x\) (add (-1.5x-1.75) to both sides) \(\large \Rightarrow x=-5\)